OpenStudy (anonymous):

-5x + 4(2x + 2) = -1 distribute then solve for x

OpenStudy (diyadiya):

do you know to do 4(2x+2) ?

OpenStudy (anonymous):

8x+8

OpenStudy (anonymous):

?

OpenStudy (diyadiya):

right :)

OpenStudy (diyadiya):

-5x + 4(2x + 2) = -1 -5x + 8x +8 =-1

OpenStudy (diyadiya):

got this part?

OpenStudy (anonymous):

3x+8=-1

OpenStudy (anonymous):

yes

OpenStudy (diyadiya):

Righttt there :) now what will be the next step?

OpenStudy (anonymous):

3x=-9

OpenStudy (anonymous):

-3

OpenStudy (diyadiya):

Good job! :)

OpenStudy (anonymous):

x=-3(:

OpenStudy (diyadiya):

right!:D